A car traveled 120 miles on paved highways before traveling 40 miles on dirt roads. The car went three times faster on the paved roads than on the dirt roads. The total time spent traveling was four hours. What was the Car's speed on the paved roads and on the dirt roads?
Answer by Alan3354(69443)

A car traveled 120 miles on paved highways before traveling 40 miles on dirt roads. The car went three times faster on the paved roads than on the dirt roads. The total time spent traveling was four hours. What was the Car's speed on the paved roads and on the dirt roads?
----------------
3 times faster is 4 times as fast.
----
120/4r + 40/r = 4
30 + 40 = 4r
r = 70/4 = 17.5 mi/hr on unpaved
--> 70 mi/hr on paved road
--------------------------------
PS Some think "3 times faster" is 3 times as fast. If so, then what is 1 time faster, or 100% faster? If they're equal, why is it called faster? I had an argument with 2 NASA guys about that, they lost.
---------------
These are the same people who say "same difference" when talking of only 2 objects, or "we'll be there momentarily" when they mean soon.